Here is a table showing fictitious trial results of a new face cream.

cream placebo
skin improved 800 600
No improvement 400 200

a) The sample size (total number of people in the study) is:

b) The estimated probability that someone's skin improved (regardless of which skin cream was used) is:

c) The experimental probability that someone's skin improved, given that they used the new cream, is:

d) The experimental probability that someone's skin improved, given that they used the placebo, is:

e) The experimental probability that someone used the new cream, given their skin improved is:

f) The experimental probability that someone used the new cream, given their skin did not improve is:
Answer #1

(a) Sample size = 800 + 400 + 600 + 200 = 2000

(b) Probability = 1400/2000 = 0.7

(c) Probability = 800/1200 = 0.67

(d) Probability = 600/800 = 0.75

(e) Probability = 800/1400 = 0.57

(f) Probability = 400/1200 = 0.33

Please give me a thumbs-up if this helps you out. Thank you!